# The (R)Evolution of Multimodal Large Language Models: A Survey

### v1

@@ Abstract
Connecting text and visual modalities plays an essential role in generative intelligence. For this reason, inspired by the success of large language models, significant research efforts are being devoted to the development of Multimodal Large Language Models (MLLMs). These models can seamlessly integrate visual and textual modalities, both as input and output, while providing a dialogue-based interface and instruction-following capabilities. In this paper, we provide a comprehensive review of recent visual-based MLLMs, analyzing their architectural choices, multimodal alignment strategies, and training techniques. We also conduct a detailed analysis of these models across a wide range of tasks, including visual grounding, image generation and editing, visual understanding, and domain-specific applications. Additionally, we compile and describe training datasets and evaluation benchmarks, conducting comparisons among existing models in terms of performance and computational requirements. Overall, this survey offers a comprehensive overview of the current state of the art, laying the groundwork for future MLLMs.
The (R)Evolution of Multimodal Large Language Models: A Survey
Davide Caffagni1111Equal contributions., Federico Cocchi1,2111Equal contributions., Luca Barsellotti1111Equal contributions., Nicholas Moratelli1111Equal contributions., Sara Sarto1111Equal contributions., Lorenzo Baraldi2111Equal contributions., Lorenzo Baraldi1, Marcella Cornia1, and Rita Cucchiara1,3 1University of Modena and Reggio Emilia, Italy 2University of Pisa, Italy 3IIT-CNR, Italy 1{name.surname}@unimore.it2{name.surname}@phd.unipi.it
@@ 1 Introduction
The introduction of the attention operator and the Transformer architecture Vaswani et al. (2017) has enabled the creation of models capable of handling various modalities on an increasingly large scale. This advancement is largely attributed to the versatility of the operator and the adaptability of the architecture. Initially, this breakthrough was leveraged for language-specific models Devlin et al. (2018); Brown et al. (2020) but quickly extended to support visual processing backbones Dosovitskiy et al. (2021) and eventually for models that integrate diverse modalities Radford et al. (2021).
@@
In a context marked by the rapid release of new models, our goal is to offer an exhaustive overview of the MLLM landscape, with a focus on models exploiting the visual modality. This overview serves as both an update on the current state and a source of inspiration for future developments. We identify three core aspects that define these models: their architecture, training methodologies, and the tasks they are designed to perform. We begin by detailing the prevalent choices for vision encoders and adapter modules that equip LLMs with cross-modal capabilities. Following this, we delve into the training processes and data utilized. We then explore the range of tasks addressed by MLLMs. The review concludes with a discussion of the persisting challenges in the field and the promising directions for future research. Further details on training data, evaluation datasets, and performance and computational requirements are reported in the supplementary material.
@@ 2.1 Preliminaries
Large Language Models. Brown et al. (2020) discovered that in-context learning, i.e., prepending the prompt with a few examples demonstrating the desired output of an LLM Chowdhery et al. (2023); Hoffmann et al. (2022); Tay et al. (2022), improves its performance, especially over unseen tasks. Generalization can be further enhanced by providing the LLM with the natural language description of the desired task for each training sample. This technique, called instruction-tuning Chung et al. (2022); Wang et al. (2022b, a); Jiang et al. (2024), turns out to be critical for aligning the behavior of an LLM with that of humans and currently empowers the most advanced LLMs, eventually boosted via reinforcement learning from human feedback (RLHF) Ouyang et al. (2022); Achiam et al. (2023); Chen et al. (2023j); Bai et al. (2023a).
PEFT. When a pre-trained LLM needs to be adapted to a specific domain or application, parameter-efficient fine-tuning (PEFT) schemes represent an important alternative to train the entire LLM, since these strategies only introduce a few new parameters. Among these, prompt-tuning Hambardzumyan et al. (2021); Lester et al. (2021); Li and Liang (2021); Liu et al. (2023j) learns a small set of vectors to be fed to the model as soft prompts before the input text. Differently, LoRA Hu et al. (2021) constrains the number of new weights by learning low-rank matrices. This technique is orthogonal to quantization methods such as QLoRA (Dettmers et al., 2023), which further decreases the memory footprint of the LLM compared to the usual half-precision weights.
@@
BLIP-2 Li et al. (2023f) ${}^{\bigstar}$ EVA ViT-g Q-Former Visual Dialogue, VQA, Captioning, Retrieval
FROMAGe Koh et al. (2023b) ${}^{\bigstar}$ CLIP ViT-L Linear Visual Dialogue, Captioning, Retrieval
@@
Kosmos-1 Huang et al. (2023a) ${}^{\lozenge}$ CLIP ViT-L Q-Former Visual Dialogue, VQA, Captioning
LLaMA-Adapter V2 Gao et al. (2023) ${}^{\blacktriangle}$ CLIP ViT-L Linear VQA, Captioning
OpenFlamingo Awadalla et al. (2023) ${}^{\bigstar}$ CLIP ViT-L XAttn LLM VQA, Captioning
@@
Flamingo Alayrac et al. (2022) ${}^{\bigstar}$ NFNet-F6 XAttn LLM Visual Dialogue, VQA, Captioning
PaLI Chen et al. (2023i) ${}^{\blacklozenge}$ ViT-e XAttn LLM Multilingual, VQA, Captioning, Retrieval
PaLI-X Chen et al. (2023g) ${}^{\blacklozenge}$ ViT-22B XAttn LLM Multilingual, VQA, Captioning
LLaVA Liu et al. (2023e) ${}^{\blacklozenge}$ CLIP ViT-L Linear Visual Dialogue, VQA, Captioning
MiniGPT-4 Zhu et al. (2023a) ${}^{\bigstar}$ EVA ViT-g Linear VQA, Captioning
mPLUG-Owl Ye et al. (2023c) ${}^{\blacktriangle}$ CLIP ViT-L Q-Former Visual Dialogue, VQA
InstructBLIP Dai et al. (2023) ${}^{\bigstar}$ EVA ViT-g Q-Former Visual Dialogue, VQA, Captioning
@@
MultiModal-GPT Gong et al. (2023) ${}^{\blacktriangle}$ CLIP ViT-L XAttn LLM Visual Dialogue, VQA, Captioning
LaVIN Luo et al. (2023) ${}^{\blacktriangle}$ CLIP ViT-L MLP Visual Dialogue, VQA, Captioning
Otter Li et al. (2023a) ${}^{\bigstar}$ CLIP ViT-L XAttn LLM VQA, Captioning
@@
Kosmos-2 Peng et al. (2023) ${}^{\lozenge}$ CLIP ViT-L Q-Former Visual Dialogue, VQA, Captioning, Referring, REC
@@
Shikra Chen et al. (2023f) ${}^{\blacklozenge}$ CLIP ViT-L Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
@@
Clever Flamingo Chen et al. (2023b) ${}^{\blacktriangle}$ CLIP ViT-L XAttn LLM Visual Dialogue, VQA, Captioning
SVIT Zhao et al. (2023a) ${}^{\blacklozenge}$ CLIP ViT-L MLP Visual Dialogue, VQA, Captioning
BLIVA Hu et al. (2024) ${}^{\bigstar}$ EVA ViT-g Q-Former+Linear Visual Dialogue, VQA, Captioning
@@
IDEFICS Laurençon et al. (2023) ${}^{\bigstar}$ OpenCLIP ViT-H XAttn LLM Visual Dialogue, VQA, Captioning
@@
Qwen-VL Bai et al. (2023b) ${}^{\blacklozenge}$ OpenCLIP ViT-bigG Q-Former Visual Dialogue, Multilingual, VQA, Captioning, REC
StableLLaVA Li et al. (2023h) ${}^{\blacklozenge}$ CLIP ViT-L Linear Visual Dialogue, VQA, Captioning
@@
Ferret You et al. (2023) ${}^{\blacklozenge}$ CLIP ViT-L Linear Visual Dialogue, Captioning, Referring, REC, GroundCap
LLaVA-1.5 Liu et al. (2023d) ${}^{\blacklozenge}$ CLIP ViT-L MLP Visual Dialogue, VQA, Captioning
@@
MiniGPT-v2 Chen et al. (2023e) ${}^{\blacktriangle}$ EVA ViT-g Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
@@
Pink Xuan et al. (2023) ${}^{\blacktriangle}$ CLIP ViT-L Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
CogVLM Wang et al. (2023c) ${}^{\blacklozenge}$ EVA ViT-E MLP Visual Dialogue, VQA, Captioning, REC
DRESS Chen et al. (2023j) ${}^{\blacktriangle}$ EVA ViT-g Linear Visual Dialogue, VQA, Captioning
LION Chen et al. (2023d) ${}^{\bigstar}$ EVA ViT-g Q-Former+MLP Visual Dialogue, VQA, Captioning, REC
mPLUG-Owl2 Ye et al. (2023d) ${}^{\blacklozenge}$ CLIP ViT-L Q-Former Visual Dialogue, VQA, Captioning
@@
SPHINX Lin et al. (2023b) ${}^{\blacklozenge}$ Mixture Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
@@
Honeybee Cha et al. (2023) ${}^{\blacklozenge}$ CLIP ViT-L ResNet blocks Visual Dialogue, VQA, Captioning
VILA Lin et al. (2023a) ${}^{\blacklozenge}$ CLIP ViT-L Linear Visual Dialogue, VQA, Captioning
@@
SPHINX-X Gao et al. (2024) ${}^{\blacklozenge}$ Mixture Linear Visual Dialogue, Multilingual, VQA, Captioning, Referring, REC
@@ 2.2 Visual Encoder
In MLLMs, one of the key components is a visual encoder, which is specifically designed to provide the LLM with the visual extracted features. It is common to employ a frozen pre-trained visual encoder while training only a learnable interface that connects visual features with the underlying LLM. The most often employed visual encoders are based on pre-trained Vision Transformer (ViT) models with a CLIP-based objective to exploit the inherent alignment of CLIP embeddings. Popular choices are the ViT-L model from CLIP Radford et al. (2021), the ViT-H backbone from OpenCLIP Wortsman et al. (2022), and the ViT-g version from EVA-CLIP Fang et al. (2023). CLIP and OpenCLIP encoders are trained on images collected from the web, employing a contrastive approach to align the correct image-text pairs. Instead, EVA-CLIP is a family of models offering a practical and effective solution for training CLIP models. In particular, EVA models are pre-trained to reconstruct the masked-out image-text aligned visual features, conditioned on visible image patches.
As shown in Li et al. (2023f), a stronger image encoder leads to better performance. Building on this insight, Lin et al. (2023b) and Gao et al. (2024) propose an ensemble of frozen visual backbones to capture robust visual representations and different levels of information granularity. Concurrently, PaLI models Chen et al. (2023i, g), noticing an imbalance between language and visual parameters, propose scaling the visual backbone respectively to a 4- and 22-billion parameter ViT.
The utilization of such large and powerful models is made feasible by the common practice of maintaining the visual encoder frozen during training, as observed in Li et al. (2023f); Huang et al. (2023a); Gao et al. (2023); Chen et al. (2023f). However, employing a frozen visual encoder has some limitations, primarily due to the constrained number of parameters, resulting in an inadequate alignment between the visual and language modalities. Specifically, the dense features, extracted from the visual model, may fragment the fine-grained image information and bring large computation due to the lengthy sequence when fed into the language model. To mitigate this issue, other approaches Ye et al. (2023c, d) employ a two-stage training paradigm. In the first stage, they incorporate a trainable visual backbone while maintaining the pre-trained LLM frozen. According to their findings, enabling the vision encoder to be trainable enhances performance on tasks such as visual question answering or visual description. However, it may lead to performance degradation in other tasks, indicating a degree of forgetting and damage to the general visual representation.

### v2

# The Revolution of Multimodal Large Language Models: A Survey

@@ Abstract
Connecting text and visual modalities plays an essential role in generative intelligence. For this reason, inspired by the success of large language models, significant research efforts are being devoted to the development of Multimodal Large Language Models (MLLMs). These models can seamlessly integrate visual and textual modalities, while providing a dialogue-based interface and instruction-following capabilities. In this paper, we provide a comprehensive review of recent visual-based MLLMs, analyzing their architectural choices, multimodal alignment strategies, and training techniques. We also conduct a detailed analysis of these models across a wide range of tasks, including visual grounding, image generation and editing, visual understanding, and domain-specific applications. Additionally, we compile and describe training datasets and evaluation benchmarks, conducting comparisons among existing models in terms of performance and computational requirements. Overall, this survey offers a comprehensive overview of the current state of the art, laying the groundwork for future MLLMs.
The Revolution of Multimodal Large Language Models: A Survey
Davide Caffagni1111Equal contribution., Federico Cocchi1,2111Equal contribution., Luca Barsellotti1111Equal contribution., Nicholas Moratelli1111Equal contribution., Sara Sarto1111Equal contribution., Lorenzo Baraldi2111Equal contribution., Lorenzo Baraldi1, Marcella Cornia1,  and Rita Cucchiara1,3 1University of Modena and Reggio Emilia, Italy 2University of Pisa, Italy 3IIT-CNR, Italy 1{name.surname}@unimore.it2{name.surname}@phd.unipi.it
@@ 1 Introduction
The introduction of the attention operation and the Transformer architecture Vaswani et al. (2017) has enabled the creation of models capable of handling various modalities on an increasingly large scale. This advancement is largely attributed to the versatility of the operator and the adaptability of the architecture. Initially, this breakthrough was leveraged for language-specific models Devlin et al. (2018); Brown et al. (2020) but quickly extended to support diverse modalities Li et al. (2019); Lu et al. (2019) and facilitate their integration within unified embedding spaces Radford et al. (2021).
@@
In a context marked by the rapid release of new models, our goal is to offer an exhaustive overview of the MLLM landscape, with a focus on models exploiting the visual modality. This overview serves as both an update on the current state and a source of inspiration for future developments. We identify three core aspects that define these models: their architecture, training methodologies, and the tasks they are designed to perform. We begin by detailing the prevalent choices for vision encoders and adapter modules that equip LLMs with cross-modal capabilities. Following this, we delve into the training processes and data utilized. We then explore the range of tasks addressed by MLLMs. The review concludes with a discussion of the persisting challenges in the field and the promising directions for future research. Further details on training data, evaluation datasets, performance and computational requirements are reported in the supplementary material.
The motivation behind this survey stems from an emerging scientific interest in the field of MLLMs, as evidenced by the constant increase in published works. In comparison with existing surveys on the topic Yin et al. (2023a); Wu et al. (2023b); Huang et al. (2023a), our paper exhibits substantial differences. Notably, it addresses several critical areas that were overlooked in prior works, including visual grounding, image generation, and editing. Furthermore, our survey details the main components utilized by each discussed MLLM, such as the visual encoders and the specific LLM employed. Additionally, our analysis offers a comparative perspective on the performance and hardware requirements of the discussed papers, incorporating both quantitative results and detailed information on benchmarks. Through this comprehensive approach, our survey aims to fill the existing gaps and provide a more nuanced understanding of the current landscape in the field.
@@ 2.1 Preliminaries
Large Language Models. Brown et al. (2020) discovered that in-context learning, i.e., prepending the prompt with a few examples demonstrating the desired output of an LLM Chowdhery et al. (2023); Hoffmann et al. (2022); Tay et al. (2022), improves its performance, especially over unseen tasks. Generalization can be further enhanced by providing the LLM with the natural language description of the desired task for each training sample. This technique, called instruction-tuning Chung et al. (2022); Wang et al. (2022b, a); Jiang et al. (2024), turns out to be critical for aligning the behavior of an LLM with that of humans and currently empowers the most advanced LLMs, eventually boosted via reinforcement learning from human feedback (RLHF) Ouyang et al. (2022); Achiam et al. (2023); Chen et al. (2023l); Bai et al. (2023a).
PEFT. When a pre-trained LLM needs to be adapted to a specific domain or application, parameter-efficient fine-tuning (PEFT) schemes represent an important alternative to train the entire LLM, since these strategies only introduce a few new parameters. Among these, prompt-tuning Hambardzumyan et al. (2021); Lester et al. (2021); Li and Liang (2021); Liu et al. (2023j) learns a small set of vectors to be fed to the model as soft prompts before the input text. Differently, LoRA Hu et al. (2021) constrains the number of new weights by learning low-rank matrices. This technique is orthogonal to quantization methods such as QLoRA (Dettmers et al., 2024), which further decreases the memory footprint of the LLM compared to the usual half-precision weights.
@@
Pre-Training of Model Components. The main components of MLLMs are the visual encoder and the language model. The visual encoder is designed to provide LLMs with visual information and the most used ones are CLIP-based architectures Radford et al. (2021); Wortsman et al. (2022) whose pre-training objective is the alignment between CLIP embeddings, obtained thanks to a contrastive loss that aligns the correct image-text pairs. An exception is the EVA-CLIP models family Fang et al. (2023), which exploits a MAE pre-training strategy He et al. (2022) to reconstruct the masked-out image-text aligned visual features, conditioned on visible image patches. On the other hand, LLMs primarily rely on the widely employed Transformer model, although the Mamba architecture Gu and Dao (2023) has also emerged in recent times. This proposes to make a State-Space Model (SSM) time-dependent, effectively creating a selective SSM with favorable properties: (i) inference costs and memory requirements that scale linearly with the sequence length, and (ii) efficient parallel training thanks to a smart GPU implementation of the algorithm. Similar to Transformers, Mamba models for language modeling are pre-trained using the next token prediction task. Very recent studies propose MLLMs featuring Mamba as the language backbone Qiao et al. (2024); Zhao et al. (2024).
@@
BLIP-2 Li et al. (2023g) FlanT5-XXL-11B EVA ViT-g Q-Former Visual Dialogue, VQA, Captioning, Retrieval
FROMAGe Koh et al. (2023b) OPT-6.7B CLIP ViT-L Linear Visual Dialogue, Captioning, Retrieval
@@
Kosmos-1 Huang et al. (2023b) Magneto-1.3B CLIP ViT-L Q-Former Visual Dialogue, VQA, Captioning
LLaMA-Adapter V2 Gao et al. (2023) LLaMA-7B CLIP ViT-L Linear VQA, Captioning
OpenFlamingo Awadalla et al. (2023) MPT-7B CLIP ViT-L XAttn LLM VQA, Captioning
@@
Flamingo Alayrac et al. (2022) Chinchilla-70B NFNet-F6 XAttn LLM Visual Dialogue, VQA, Captioning
PaLI Chen et al. (2023j) mT5-XXL-13B ViT-e XAttn LLM Multilingual, VQA, Captioning, Retrieval
PaLI-X Chen et al. (2023h) UL2-32B ViT-22B XAttn LLM Multilingual, VQA, Captioning
LLaVA Liu et al. (2023e) Vicuna-13B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning
MiniGPT-4 Zhu et al. (2023a) Vicuna-13B EVA ViT-g Linear VQA, Captioning
mPLUG-Owl Ye et al. (2023c) LLaMA-7B CLIP ViT-L Q-Former Visual Dialogue, VQA
InstructBLIP Dai et al. (2023) Vicuna-13B EVA ViT-g Q-Former Visual Dialogue, VQA, Captioning
@@
MultiModal-GPT Gong et al. (2023) LLaMA-7B CLIP ViT-L XAttn LLM Visual Dialogue, VQA, Captioning
LaVIN Luo et al. (2023) LLaMA-13B CLIP ViT-L MLP Visual Dialogue, VQA, Captioning
Otter Li et al. (2023b) LLaMA-7B CLIP ViT-L XAttn LLM VQA, Captioning
@@
Kosmos-2 Peng et al. (2023) Magneto-1.3B CLIP ViT-L Q-Former Visual Dialogue, VQA, Captioning, Referring, REC
@@
Shikra Chen et al. (2023f) Vicuna-13B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
@@
Clever Flamingo Chen et al. (2023b) LLaMA-7B CLIP ViT-L XAttn LLM Visual Dialogue, VQA, Captioning
SVIT Zhao et al. (2023a) Vicuna-13B CLIP ViT-L MLP Visual Dialogue, VQA, Captioning
BLIVA Hu et al. (2024) Vicuna-7B EVA ViT-g Q-Former+Linear Visual Dialogue, VQA, Captioning
@@
IDEFICS Laurençon et al. (2024) LLaMA-65B OpenCLIP ViT-H XAttn LLM Visual Dialogue, VQA, Captioning
@@
Qwen-VL Bai et al. (2023b) Qwen-7B OpenCLIP ViT-bigG Q-Former Visual Dialogue, Multilingual, VQA, Captioning, REC
StableLLaVA Li et al. (2023i) Vicuna-13B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning
@@
Ferret You et al. (2023) Vicuna-13B CLIP ViT-L Linear Visual Dialogue, Captioning, Referring, REC, GroundCap
LLaVA-1.5 Liu et al. (2023d) Vicuna-13B CLIP ViT-L MLP Visual Dialogue, VQA, Captioning
@@
MiniGPT-v2 Chen et al. (2023e) LLaMA-2-7B EVA ViT-g Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
@@
Pink Xuan et al. (2023) Vicuna-7B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
CogVLM Wang et al. (2023c) Vicuna-7B EVA ViT-E MLP Visual Dialogue, VQA, Captioning, REC
DRESS Chen et al. (2023l) Vicuna-13B EVA ViT-g Linear Visual Dialogue, VQA, Captioning
LION Chen et al. (2023d) FlanT5-XXL-11B EVA ViT-g Q-Former+MLP Visual Dialogue, VQA, Captioning, REC
mPLUG-Owl2 Ye et al. (2023d) LLaMA-2-7B CLIP ViT-L Q-Former Visual Dialogue, VQA, Captioning
@@
SPHINX Lin et al. (2023b) LLaMA-2-13B Mixture Linear Visual Dialogue, VQA, Captioning, Referring, REC, GroundCap
@@
Honeybee Cha et al. (2023) Vicuna-13B CLIP ViT-L ResNet blocks Visual Dialogue, VQA, Captioning
VILA Lin et al. (2023a) LLaMA-2-13B CLIP ViT-L Linear Visual Dialogue, VQA, Captioning
@@
SPHINX-X Gao et al. (2024) Mixtral-8$\times$7B Mixture Linear Visual Dialogue, Multilingual, VQA, Captioning, Referring, REC
@@ 2.2 Visual Encoder
In MLLMs, one of the key components is a visual encoder, which is specifically designed to provide the LLM with the visual extracted features. It is common to employ a frozen pre-trained visual encoder while training only a learnable interface that connects visual features with the underlying LLM. While this is usually done using low-resolution images with fixed aspect ratios, some attempts Xu et al. (2024); Li et al. (2023l) involve adapting pre-trained visual backbones to handle images of different resolutions and aspect ratios. Further details on how to handle higher-resolution images are provided in the supplementary.
The most often employed visual encoders are based on pre-trained Vision Transformer (ViT) models with a CLIP-based objective to exploit the inherent alignment of CLIP embeddings. Popular choices are the ViT-L model from CLIP Radford et al. (2021), the ViT-H backbone from OpenCLIP Wortsman et al. (2022), and the ViT-g version from EVA-CLIP Fang et al. (2023).
As shown in Li et al. (2023g), a stronger image encoder leads to better performance. Building on this insight, Lin et al. (2023b) and Gao et al. (2024) propose an ensemble of frozen visual backbones to capture robust visual representations and different levels of information granularity. Concurrently, PaLI models Chen et al. (2023j, h), noticing an imbalance between language and visual parameters, propose scaling the visual backbone respectively to a 4- and 22-billion parameter ViT.
The utilization of such large and powerful models is made feasible by the common practice of maintaining the visual encoder frozen during training, as observed in Li et al. (2023g); Huang et al. (2023b); Gao et al. (2023); Chen et al. (2023f). However, employing a frozen visual encoder has some limitations, primarily due to the constrained number of parameters, resulting in an inadequate alignment between the visual and language modalities. Specifically, the dense features, extracted from the visual model, may fragment the fine-grained image information and bring large computation due to the lengthy sequence when fed into the language model. To mitigate this issue, other approaches Ye et al. (2023c, d) employ a two-stage training paradigm. In the first stage, they incorporate a trainable visual backbone while maintaining the pre-trained LLM frozen. According to their findings, enabling the vision encoder to be trainable enhances performance on tasks such as visual question answering or visual description. However, it may lead to performance degradation in other tasks, indicating a degree of forgetting and damage to the general visual representation.